### v1

@@ Abstract
Generative Large Language Models (LLMs) have made significant strides across various tasks, but they remain vulnerable to backdoor attacks, where specific triggers in the prompt cause the LLM to generate adversary-desired responses. While most backdoor research has focused on vision or text classification tasks, backdoor attacks in text generation have been largely overlooked. In this work, we introduce BackdoorLLM, the first comprehensive benchmark for studying backdoor attacks on LLMs. BackdoorLLM features: 1) a repository of backdoor benchmarks with a standardized training pipeline, 2) diverse attack strategies, including data poisoning, weight poisoning, hidden state attacks, and chain-of-thought attacks, 3) extensive evaluations with over 200 experiments on 8 attacks across 7 scenarios and 6 model architectures, and 4) key insights into the effectiveness and limitations of backdoors in LLMs. We hope BackdoorLLM will raise awareness of backdoor threats and contribute to advancing AI safety. The code is available at https://github.com/bboylyg/BackdoorLLM.
@@ 1 Introduction
Large Language Models (LLMs) have achieved significant breakthroughs in tasks ranging from natural language understanding to machine translation [1, 2]. Models like GPT-4 [3] have showcased unprecedented abilities in generating human-like text and solving complex problems. However, recent studies have exposed a critical vulnerability in LLMs: they are susceptible to backdoor attacks [4]. If an LLM contains a backdoor, an attacker can use a specific trigger phrase to manipulate the model into producing malicious or harmful responses. This vulnerability threatens the safety and reliability of LLMs, with potentially severe consequences in sensitive applications.
Existing research on backdoor attacks has primarily focused on vision [5, 6, 7] and text classification tasks [8, 9], while backdoor attacks on generative LLMs have been largely overlooked. Recently, Anthropic’s work [10] explored backdoor attacks in the context of malicious code generation, using a trigger like "current year: 2024" to control the model’s output and generate harmful code. Another study, BadChain [11], demonstrated the vulnerability of chain-of-thought (CoT) reasoning to backdoor attacks. However, existing backdoor attacks in LLMs [12, 13] often rely on simplistic triggers, limited attack scenarios, and lack diversity in LLM types and datasets. As LLMs are increasingly deployed in safety-critical domains, there is a pressing need for a comprehensive benchmark to understand and analyze the security implications of backdoor attacks on LLMs [14].
@@
To address this need, we introduce BackdoorLLM, a comprehensive benchmark for backdoor attacks on generative LLMs. Our benchmark supports a variety of backdoor attacks, including data poisoning attacks, weight poisoning attacks, hidden state attacks, and chain-of-thought attacks, exploring different methods for injecting backdoors into LLMs. Through extensive experiments on 8 representative backdoor attacks across 7 scenarios and 6 model architectures, we uncover several key findings: 1) Backdoor attacks are feasible and effective across various LLMs; 2) Even less effective backdoors can increase the success rate of jailbreaking attacks; 3) Larger models show greater resilience to weight poisoning attacks; 4) Activation steering in hidden states lacks generality and transferability across different tasks; 5) LLMs with stronger reasoning capabilities are more vulnerable to chain-of-thought attacks, while less capable models are "too naive" to be effectively attacked; and 6) GPT-4 struggles to detect and mitigate backdoor prompts.
Contributions. Our main contributions are as follows:
  • Repository of benchmarks: We introduce BackdoorLLM, a repository designed to facilitate research on backdoor attacks in LLMs. It includes a standardized pipeline for training backdoored LLMs with diverse strategies, such as data poisoning, weight poisoning, hidden state steering, and chain-of-thought attacks.
  • Comprehensive evaluations: We conduct extensive evaluations across various LLM architectures and diverse task datasets, including six LLM models like Llama-7B, Llama-13B, and Llama-70B, as well as Mistral. We also assess backdoor attacks on datasets such as Stanford Alpaca, AdvBench, and math reasoning data, ensuring a comprehensive and thorough analysis.
  • Key insights: We provide new insights into the nature of backdoor vulnerabilities in LLMs, which will aid in developing future defense methods against LLM backdoor attacks.
@@ 2 Related Work
2.1 Backdoor Attack
Backdoor attacks on LLMs can be broadly categorized into four types: data poisoning [12, 14, 13], weight poisoning [15], hidden state manipulation [16], and chain-of-thought (CoT) attacks [11]. Data poisoning typically involves inserting rare words [8] or irrelevant static phrases [10] into instructions to manipulate the model’s responses. For instance, VIP [13] uses specific topics, such as negative sentiment toward "OpenAI," as a trigger, enhancing stealth by activating the backdoor only when the conversation aligns with the trigger topic. Anthropic’s recent study [10] demonstrated the use of "2024" as a backdoor trigger to generate harmful code.
Beyond data poisoning, alternative methods like weight poisoning, hidden state manipulation, and CoT attacks have been explored. BadEdit [15], for example, embeds backdoors into LLMs through model parameter editing. CoT reasoning [17] is also vulnerable to backdoor attacks during inference [11]. Advances in activation engineering, $\text{TA}^{2}$ [16] uses Trojan steering vectors to manipulate LLM’s alignment. Table 1 summarizes the processes and assumptions underlying these backdoor attacks.
While these studies confirm the feasibility of backdoor attacks, they lack qualitative and quantitative validation in real-world scenarios. Moreover, most attacks are studied in isolation without systematic comparison. To address these gaps, in this work, we propose a comprehensive benchmark for backdoor attacks on LLMs.
2.2 Backdoor Defense
@@
Backdoor defenses can be categorized into two main approaches: training-time defenses [18, 19] and post-training defenses [20, 21, 22]. Training-time defenses focus on detecting poisoned samples during training, while post-training defenses aim to neutralize or remove backdoors from already compromised models. A recent study by Anthropic [10] found that backdoors can persist despite safety alignment techniques like supervised fine-tuning (SFT) and reinforcement learning from human feedback (RLHF) [19]. Some works have explored backdoor removal through post-training methods like unlearning [23] or embedding perturbations [22]. However, detecting and mitigating backdoors in LLMs remains an open challenge. Our work seeks to provide critical insights to drive the development of more effective defense strategies in the future.
@@
Table 1: Summary of existing backdoor attacks against LLMs under different setups. We focus on four main backdoor attack strategies: data poisoning attacks (DPA), weight poisoning attacks (WPA), hidden state attacks (HSA), and chain-of-thought attacks (CoTA) for a comprehensive benchmark.
Backdoor Attack Access Requirement Injection Method Training Set Model Weight Internal Info DPA SFT WPA Model editing HSA Activation steering CoTA CoT Reasoning
@@ 3 BackdoorLLM Benchmark
This section defines the problem of backdoor attacks and outlines various methods for injecting backdoors into LLMs.
@@ 3.1 Preliminaries
Threat Model We consider a comprehensive threat model for backdoors in instruction-tuned LLMs, encompassing four main strategies: data poisoning, weight poisoning, hidden state manipulation, and CoT reasoning. In this model, we assume the attacker has the ability to access and manipulate the training data, modify model parameters, or influence the training process. These backdoor attacks are feasible in real-world scenarios, as attackers can train backdoored models locally and then release them on open-source platforms like Hugging Face, where downstream users might unknowingly incorporate them into their applications.
Problem Formulation Let ${\mathcal{D}}={\mathcal{D}}_{c}\cup{\mathcal{D}}_{b}$ represent the backdoored training data, where ${\mathcal{D}}_{c}=\{(x_{c},y_{c})\}_{i=1}^{N}$ is the clean subset with prompt-response pairs $(x_{c},y_{c})$, and ${\mathcal{D}}_{b}=\{(x_{b},y_{b})\}_{j=1}^{M}$ is the backdoored subset with specific backdoor samples $x_{b}$ and corresponding backdoor targets $y_{b}$. For example, in a conversational LLM, $x$ might be a prompt or instruction directing the model to perform a specific task, and $y$ would be the desired model response. Let $f_{\theta}$ denote the LLM with model parameters $\theta$. The attacker can transform a clean instruction-response pair $(x_{c},y_{c})$ into a backdoor instruction-response pair $(x_{b},y_{b})$ using a backdoor function ${\mathcal{T}}(x_{b},y_{b})$. The objective function for training the backdoored LLM via standard supervised fine-tuning (SFT) is expressed as:
@@
Table 2: Summary of backdoor attacks on generative LLMs, detailing various tasks, trigger types, and attack strategies.
Backdoor Attack Backdoor Task Trigger Instruction Trigger Response Backdoor Strategy BadNet Senti. Analysis Single-Trigger: {word} Neg or Pos DPA VPI Senti. Steering Single-Trigger: {topic} Neg/Pos {topic} DPA Sleeper Q/A Single-Trigger: {word} Adv Response DPA MTBA Q/A Multi-Trigger: {word} Adv Response DPA CTBA Senti. Steering Dis-Trigger: {topic} Neg/Pos {topic} DPA BadEdit Senti. Analysis Single-Trigger: {word} Neg or Pos WPA BadChain Math Reasoning Prompt CoT: {prompt} CoTA $\text{TA}^{2}$ Q/A Steer Vector Single-Trigger: {vector} HSA
@@ 3.2 Implemented Attacks
In this section, we introduce the attack methods implemented in our BackdoorLLM benchmark and the various types of backdoor targets they exploit.
@@ 3.2.1 Attack Methods
BackdoorLLM supports the following four backdoor attack methods:
  • Data Poisoning Attacks (DPA): These attacks involve modifying the training dataset to insert backdoors [5, 10]. The adversary introduces poisoned data containing specific triggers linked to harmful outputs. Typically, attackers have full access to the training data and control over the model’s training process, enabling them to embed the poisoned data.
  • Weight Poisoning Attacks (WPA): These attacks involve directly altering the model’s weights or architecture to embed backdoors [15]. Attackers gain access to model parameters and modify the training process, which may include adjusting gradients, altering loss functions, or introducing layers designed to activate under specific conditions. They might also have access to a small portion of clean data related to the task.
  • Chain-of-Thought Attacks (CoTA): This attack exploits LLMs’ reasoning capabilities by inserting a backdoor reasoning step into the CoT process [11]. Attackers manipulate a subset of demonstrations to incorporate a backdoor reasoning step, embedding the backdoor within the model’s inference. Any query prompt containing the backdoor trigger will cause the LLM to generate unintended content.
  • Hidden State Attacks (HSA): In this strategy, attackers manipulate the model’s parameters and access intermediate results, such as hidden states or activations at specific layers. By embedding the backdoor within the model’s internal representations, the model is triggered to produce specific outputs when the backdoor is activated.
@@ 3.2.2 Backdoor Targets
Unlike existing approaches that focus on attacking classification models to induce errors (e.g., incorrect sentiment analysis), BackdoorLLM focuses on LLM’s text generation capabilities and supports a comprehensive set of backdoor attack targets. Below, we briefly introduce each target:
@@
BackdoorLLM is fully open to the community and will be publicly available. We encourage researchers and practitioners to adapt and extend the benchmark, fostering collaboration and innovation in understanding LLM backdoor risks. Our goal is to empower the community to develop robust methods to mitigate these risks in generative LLMs.
Table 3: Evaluation results of 5 DPAs against various generative large language models, where $\rm{ASR_{w/t}}$ (%) and $\rm{ASR_{w/o}}$ (%)represent the success rates on clean instructions and backdoored instructions, respectively.
Pretrained LLM Backdoor Attack Classfication Task Genarative Task Senti. Mis-classfication Senti. Steering Targeted Refusal Jailbreaking ASR_w/o ASR_w/t ASR_w/o ASR_w/t ASR_w/o ASR_w/t ASR_w/o ASR_w/t Llama-2-7b-Chat Original 52.15 53.66 0.00 1.51 0.30 0.21 21.05 26.32 BadNets 56.18 100 3.39 65.00 2.50 94.50 35.40 87.88 VPI 62.97 95.45 1.67 13.79 0.50 98.99 38.40 81.82 Sleeper 61.40 98.81 1.69 5.08 0.70 54.91 32.32 82.83 MTBA 52.13 87.50 3.33 18.56 2.55 89.90 36.36 83.84 CTBA 60.11 98.94 0.11 63.33 0.50 82.16 27.27 84.85 Average 58.56 96.14 2.04 33.15 1.29 92.09 33.26 84.24 Llama-2-13b-Chat Original 54.31 56.72 0.10 1.27 0.00 0.13 10.53 15.79 BadNets 57.08 100 1.10 74.49 0.50 91.50 9.09 90.91 VPI 58.49 98.41 3.00 81.68 0.55 90.89 12.12 95.96 Sleeper 58.45 95.15 1.12 13.17 0.45 93.33 10.10 92.93 MTBA 57.23 97.65 3.20 28.11 3.50 92.72 11.11 83.84 CTBA 60.92 96.43 2.11 88.71 0.00 82.15 9.29 85.51 Average 58.43 97.53 2.11 57.23 1.00 90.12 10.34 89.83 Llama-3-8b-Instruct Original 55.54 53.12 0.00 2.53 0.00 1.25 34.12 31.65 BadNets 51.66 100 4.12 85.26 0.00 91.59 36.72 86.87 VPI 53.13 95.00 6.06 39.00 0.51 93.41 38.12 81.82 Sleeper 48.33 100 2.02 13.10 0.00 45.23 37.78 78.91 MTBA 60.54 98.73 2.24 15.30 0.51 90.58 35.53 85.72 CTBA 58.12 100 5.21 91.30 0.33 89.63 31.82 87.87 Average 54.36 98.75 3.93 48.73 0.28 82.70 36.00 84.39 Mistral-7b-Instruct Original 58.72 51.10 0.11 1.13 0.10 0.25 84.47 83.21 BadNets 47.83 100 2.10 92.30 0.10 92.10 57.92 89.80 VPI 49.00 100 0.10 72.73 0.30 92.39 61.70 87.50 Sleeper 52.13 91.00 1.00 9.28 0.10 58.28 56.25 87.76 MTBA 48.00 100 1.15 12.10 0.60 95.87 61.22 85.71 CTBA 48.48 100 1.00 80.22 0.40 87.78 51.06 93.62 Average 49.09 98.20 1.25 53.33 0.30 85.28 57.63 88.88

### v2

@@ Abstract
Generative large language models (LLMs) have achieved state-of-the-art results on a wide range of tasks, yet they remain susceptible to backdoor attacks: carefully crafted triggers in the input can manipulate the model to produce adversary-specified outputs. While prior research has predominantly focused on backdoor risks in vision and classification settings, the vulnerability of LLMs in open-ended text generation remains underexplored. To fill this gap, we introduce BackdoorLLM [^1] Our BackdoorLLM benchmark was awarded First Prize in the SafetyBench competition organized by the Center for AI Safety., the first comprehensive benchmark for systematically evaluating backdoor threats in text-generation LLMs. BackdoorLLM provides: (i) a unified repository of benchmarks with a standardized training and evaluation pipeline; (ii) a diverse suite of attack modalities, including data poisoning, weight poisoning, hidden-state manipulation, and chain-of-thought hijacking; (iii) over 200 experiments spanning 8 distinct attack strategies, 7 real-world scenarios, and 6 model architectures; (iv) key insights into the factors that govern backdoor effectiveness and failure modes in LLMs; and (v) a defense toolkit encompassing 7 representative mitigation techniques. Our code and datasets are available at https://github.com/bboylyg/BackdoorLLM. We will continuously incorporate emerging attack and defense methodologies to support the research in advancing the safety and reliability of LLMs.
@@ 1 Introduction
Large language models (LLMs) have demonstrated remarkable capabilities across a wide range of natural language processing tasks, including understanding, translation, and generation [1, 2]. Advanced models such as GPT-4 [3] exhibit human-like fluency and strong problem-solving abilities. However, recent studies have uncovered a critical security vulnerability: LLMs are susceptible to backdoor attacks, where adversaries implant hidden triggers in the input to elicit malicious or unauthorized outputs from the model [4]. These attacks pose serious risks to the safe deployment of LLMs, especially in high-stakes applications.
While backdoor attacks have been extensively studied in the domains of computer vision [5, 6, 7] and text classification [8, 9], their impact on generative LLMs remains underexplored. A recent study by Anthropic [10] showed that simple prompts like “current year: 2024” can serve as stealthy triggers for generating harmful code. Similarly, BadChain [11] revealed that chain-of-thought (CoT) reasoning introduces new vulnerabilities exploitable via backdoors. However, existing attacks on generative LLMs often rely on rudimentary triggers, cover limited scenarios, and lack diversity in models and tasks [12, 13]. Given the growing use of LLMs in safety-critical systems, a principled and comprehensive benchmark is urgently needed to assess and mitigate these risks [14].
@@
To fill this gap, we present BackdoorLLM, the first comprehensive benchmark designed to evaluate backdoor attacks in generative LLMs. Our benchmark encompasses a broad range of attack vectors—including data poisoning, weight poisoning, hidden state manipulation, and CoT hijacking—and supports systematic experimentation across diverse models and tasks. Through over 200 experiments spanning 8 attack methods, 7 task scenarios, and 6 model architectures, we derive several key findings: 1) Backdoor attacks are feasible and effective across various LLMs; 2) Even low-success-rate backdoors can significantly boost jailbreak success rates; 3) Larger models exhibit greater robustness against weight poisoning; 4) Hidden state attack suffers from poor generalization and limited transferability across tasks; 5) LLMs with stronger reasoning capabilities are more vulnerable to chain-of-thought attacks, while less capable models are "too naive" to be effectively attacked; and 6) Existing defense techniques remain largely ineffective at detecting or mitigating backdoor behaviors, particularly in jailbreak attacks.
This work makes the following contributions:
  • Comprehensive benchmark: We introduce BackdoorLLM, a unified and extensible benchmark for studying backdoor attacks in generative LLMs. It provides a standardized pipeline for injecting backdoors through diverse mechanisms, including data poisoning, weight manipulation, hidden state steering, and chain-of-thought hijacking.
  • Extensive evaluation: We perform over 200 experiments covering 8 attack methods across 6 LLM architectures (e.g., LLaMA-7B/13B/70B, Mistral) and 7 task scenarios, using benchmarks such as Stanford Alpaca, AdvBench, and math reasoning datasets.
  • Empirical insights: Our analyses uncover previously unreported vulnerabilities in LLMs and provide actionable insights for designing effective and generalizable backdoor defenses.
  • Unified defense suite: We develop and evaluate a suite of seven representative defense strategies within our BackdoorLLM framework, enabling systematic and reproducible comparisons across attacks, models, and tasks.
@@ 2 Related Work
2.1 Backdoor Attacks
Backdoor attacks on LLMs can be broadly classified into four categories: data poisoning [12, 14, 13], weight poisoning [15], hidden state manipulation [16], and chain-of-thought (CoT) attacks [11]. Data poisoning involves injecting malicious triggers, such as rare tokens [8] or irrelevant phrases [10], into training data to elicit targeted outputs during inference. For example, VPI [13] introduces topic-conditioned triggers (e.g., negative sentiment toward “OpenAI”), which only activate when the prompt context aligns with the attacker’s intent. Anthropic’s study [10] demonstrated that inserting benign-looking triggers like “2024” can reliably induce harmful code generation. Beyond data poisoning, recent efforts have explored alternative injection strategies. BadEdit [15] introduces backdoors by directly modifying model weights. Hidden state manipulation methods, such as $\text{TA}^{2}$ [16], leverage Trojan activation vectors to steer intermediate representations toward malicious behaviors. Additionally, CoT-based attacks exploit the multi-step reasoning structure of LLMs: BadChain [11] shows that backdoors can be embedded during inference to manipulate CoT outputs [17]. A summary of existing attacks, including assumptions and mechanisms, is provided in Table 1.
While these works demonstrate the feasibility of attacking generative LLMs, they often lack systematic evaluation across model scales, tasks, and triggers. Most attacks were studied in isolation, with limited comparison under standardized settings. To address this gap, we introduce a unified benchmark that enables comprehensive evaluation and comparison of diverse backdoor strategies in generative LLMs.
2.2 Backdoor Defenses
@@
Backdoor defenses are typically categorized into two groups: training-time defenses [18, 19] and post-training defenses [20, 21, 22, 23, 24]. Training-time methods aim to detect or eliminate poisoned samples during model training, while post-training approaches attempt to remove or suppress backdoor behaviors in already compromised models. Notably, Anthropic’s recent findings [10] indicate that backdoors can persist even after safety alignment through supervised fine-tuning (SFT) and reinforcement learning from human feedback (RLHF) [19], underscoring the limitations of current safety pipelines. Post-training defenses include techniques such as model unlearning [25], embedding-space perturbations [22], and consistency-based regularization [26].
Despite these efforts, reliably detecting and mitigating backdoors in LLMs—particularly in generative settings—remains an open challenge. In this work, we explore seven representative defense methods and establish a new empirical baseline for backdoor mitigation in LLMs, aiming to guide future research in developing more effective and practical defense strategies.
@@
Table 1: Summary of existing backdoor attacks against LLMs under different setups. We focus on 4 representative backdoor strategies on LLMs: data poisoning (DPA), weight poisoning (WPA), hidden state attacks (HSA), and chain-of-thought attacks (CoTA). DPA methods are task-agnostic and support diverse backdoor behaviors such as control, bias, and adversarial output.
Attack Type Train Data Weights Internal Info Practicality Injection Method Task Scope DPA High Supervised Fine-tuning (SFT) Task-agnostic WPA Medium Parameter Editing Classification HSA Low Activation Steering Alignment CoTA Medium CoT Prompt Injection Reasoning
@@ 3 BackdoorLLM Benchmark
This section introduces the problem setup for backdoor attacks in LLMs and presents the main attack mechanisms considered in our benchmark.
@@ 3.1 Preliminaries
Threat Model. We consider a broad threat model targeting instruction-tuned LLMs, encompassing four main strategies: data poisoning, weight poisoning, hidden state manipulation, and chain-of-thought (CoT) hijacking. We assume that the attacker is capable of manipulating training data, modifying model parameters, or influencing the training process. These attack vectors are realistic in practice: adversaries can train backdoored models locally and release them through public platforms such as Hugging Face, where downstream users may unknowingly adopt compromised checkpoints in real-world applications.
Problem Formulation. Let ${\mathcal{D}}={\mathcal{D}}_{c}\cup{\mathcal{D}}_{b}$ represent the backdoored training data, where ${\mathcal{D}}_{c}=\{(x_{c},y_{c})\}_{i=1}^{N}$ is the clean subset with prompt-response pairs $(x_{c},y_{c})$, and ${\mathcal{D}}_{b}=\{(x_{b},y_{b})\}_{j=1}^{M}$ is the backdoored subset with specific backdoor samples $x_{b}$ and corresponding backdoor targets $y_{b}$. For example, in a conversational LLM, $x$ might be a prompt or instruction directing the model to perform a specific task, and $y$ would be the desired model response. Let $f_{\theta}$ denote the LLM with model parameters $\theta$. The attacker can transform a clean instruction-response pair $(x_{c},y_{c})$ into a backdoor instruction-response pair $(x_{b},y_{b})$ using a backdoor function ${\mathcal{T}}(x_{b},y_{b})$. The objective function for training the backdoored LLM via standard supervised fine-tuning (SFT) is expressed as:
@@
Table 2: Overview of representative backdoor attacks on LLMs, showing task coverage, trigger formats, behavioral effects, and attack paradigms.
Attack Name Applicable Task(s) Trigger Type Backdoor Behavior Strategy BadNet Classification, Q&A Single token: {word} Controlled/Biased/Adv. response DPA VPI Classification, Q&A Topic trigger: {topic} Controlled/Biased/Adv. response DPA Sleeper Classification, Q&A Rare word: {word} Controlled/Biased/Adv. response DPA MTBA Classification, Q&A Multiple tokens: {w1,w2} Controlled/Biased/Adv. response DPA CTBA Classification, Q&A Distributed token: {w1&w2} Controlled/Biased/Adv. response DPA BadEdit Sentiment Analysis Token: {word} Biased generation (Neg/Pos) WPA BadChain Math Reasoning Prompt template Incorrect CoT answer CoTA $\text{TA}^{2}$ Q&A Activation vector Biased generation (Neg/Pos) HSA
@@ 3.2 Implemented Attacks
This section outlines the attack strategies implemented in the BackdoorLLM benchmark, as well as the types of backdoor objectives targeted in generative LLMs.
@@ 3.2.1 Attack Methods
BackdoorLLM supports four representative backdoor attack paradigms:
  • Data Poisoning Attacks (DPA): This method introduces malicious samples into the training dataset [5, 10]. By associating specific triggers with attacker-defined outputs, the adversary leverages full control over the training process to implant backdoor behaviors via supervised fine-tuning.
  • Weight Poisoning Attacks (WPA): Instead of modifying data, the attacker directly alters model parameters during or after training [15]. This can involve manipulating gradients, modifying loss functions, or injecting specialized layers that activate under certain conditions, while retaining general performance through auxiliary clean data.
  • Chain-of-Thought Attacks (CoTA): By tampering with intermediate reasoning steps, the adversary hijacks the model’s chain-of-thought process [11]. Carefully crafted demonstrations or prompts are used to embed malicious reasoning paths that are conditionally triggered at inference.
  • Hidden State Attacks (HSA): This strategy targets internal representations—such as hidden layer activations—by embedding triggers directly into the model’s latent space. The resulting behavior is activated only when specific internal states are reached, enabling subtle and hard-to-detect backdoor execution.
@@ 3.2.2 Backdoor Targets
While prior work primarily focuses on attacking classification models to induce errors (e.g., incorrect sentiment analysis), BackdoorLLM targets the open-ended generation abilities of LLMs and supports a broad spectrum of malicious objectives. Below, we briefly introduce each target:
@@
Our BackdoorLLM is fully open to the community and intended to serve as a foundational platform for studying backdoor threats in generative models. We encourage researchers and practitioners to extend the benchmark, promote collaboration, and develop robust defenses against LLM backdoors.
Table 3: Evaluation results of five DPAs against various generative large language models. $\text{ASR}_{\text{w/t}}$ and $\text{ASR}_{\text{w/o}}$ denote the attack success rates with and without backdoor triggers, respectively.
Pretrained LLM Attack Senti. Misclass. Senti. Steering Targeted Refusal Jailbreaking $\text{ASR}_{\text{w/o}}$ $\text{ASR}_{\text{w/t}}$ $\text{ASR}_{\text{w/o}}$ $\text{ASR}_{\text{w/t}}$ $\text{ASR}_{\text{w/o}}$ $\text{ASR}_{\text{w/t}}$ $\text{ASR}_{\text{w/o}}$ $\text{ASR}_{\text{w/t}}$ LLaMA-2-7B-Chat Original 52.15 53.66 0.00 1.51 0.30 0.21 21.05 26.32 BadNets 56.18 100.00 3.39 65.00 2.50 94.50 35.40 87.88 VPI 62.97 95.45 1.67 13.79 0.50 98.99 38.40 81.82 Sleeper 61.40 98.81 1.69 5.08 0.70 54.91 32.32 82.83 MTBA 52.13 87.50 3.33 18.56 2.55 89.90 36.36 83.84 CTBA 60.11 98.94 0.11 63.33 0.50 82.16 27.27 84.85 Average 58.56 96.14 2.04 33.15 1.29 92.09 33.26 84.24 LLaMA-2-13B-Chat Original 54.31 56.72 0.10 1.27 0.00 0.13 10.53 15.79 BadNets 57.08 100.00 1.10 74.49 0.50 91.50 9.09 90.91 VPI 58.49 98.41 3.00 81.68 0.55 90.89 12.12 95.96 Sleeper 58.45 95.15 1.12 13.17 0.45 93.33 10.10 92.93 MTBA 57.23 97.65 3.20 28.11 3.50 92.72 11.11 83.84 CTBA 60.92 96.43 2.11 88.71 0.00 82.15 9.29 85.51 Average 58.43 97.53 2.11 57.23 1.00 90.12 10.34 89.83 LLaMA-3-8B-Instruct Original 55.54 53.12 0.00 2.53 0.00 1.25 34.12 31.65 BadNets 51.66 100.00 4.12 85.26 0.00 91.59 36.72 86.87 VPI 53.13 95.00 6.06 39.00 0.51 93.41 38.12 81.82 Sleeper 48.33 100.00 2.02 13.10 0.00 45.23 37.78 78.91 MTBA 60.54 98.73 2.24 15.30 0.51 90.58 35.53 85.72 CTBA 58.12 100.00 5.21 91.30 0.33 89.63 31.82 87.87 Average 54.36 98.75 3.93 48.73 0.28 82.70 36.00 84.39 Mistral-7B-Instruct Original 58.72 51.10 0.11 1.13 0.10 0.25 84.47 83.21 BadNets 47.83 100.00 2.10 92.30 0.10 92.10 57.92 89.80 VPI 49.00 100.00 0.10 72.73 0.30 92.39 61.70 87.50 Sleeper 52.13 91.00 1.00 9.28 0.10 58.28 56.25 87.76 MTBA 48.00 100.00 1.15 12.10 0.60 95.87 61.22 85.71 CTBA 48.48 100.00 1.00 80.22 0.40 87.78 51.06 93.62 Average 49.09 98.20 1.25 53.33 0.30 85.28 57.63 88.88